# Scaling relations for heat and momentum transport in sheared Rayleigh–Bénard convection

### v1

@@ 1 Introduction
The interplay between buoyancy and shear in mixed thermal convection can be studied by either adding Couette–type forcing to the Rayleigh–Bénard (RB) system (Ahlers et al., 2009; Lohse & Xia, 2010; Chilla & Schumacher, 2012; Xia, 2013; Shishkina, 2021; Ahlers et al., 2022; Lohse & Shishkina, 2023) to obtain the Couette–RB (CRB) system (Ingersoll, 1966; Hathaway & Somerville, 1986; Domaradzki & Metcalfe, 1988; Solomon & Gollub, 1990; Shevkar et al., 2019; Blass et al., 2020, 2021), or by applying a Poiseuille–type forcing to obtain the Poiseuille–RB (PRB) system (Zonta & Soldati, 2014; Scagliarini et al., 2014, 2015; Pirozzoli et al., 2017). A schematic of the two systems is shown in figure 1.
@@
Parameter Definition Type
Short description
$Ra$ $\bigl{(}\beta gH^{3}\Delta\bigr{)}/(\nu\kappa)$ Control
Rayleigh number
$Pr$ $\nu/\kappa$ Control
Prandtl number
$Re_{F}$ $U_{F}H/\nu\equiv\sqrt{Ra/Pr}$ Control
Reynolds number associated with free-fall velocity
$Re_{w}$ $U_{w}H/\nu$ Control
Wall Reynolds number for CRB system
$Re_{b}$ $U_{b}H/\nu$ Control
Bulk Reynolds number for PRB system
$Re_{S}$ $U_{S}H/\nu$ Control
Reynolds number associated with shear forcing
$Re_{L}$ $U_{L}H/\nu$ Response
Reynolds number associated with LSC rolls for sheared RB
$Re_{R}$ $U_{R}H/\nu\equiv Re_{L}(Re_{S}=0)$ Response
Reynolds number associated with LSC rolls for pure RB (“wind Reynolds number”)
$Re_{T}$ $U_{T}H/\nu$ Response
Reynolds number computed using the total velocity
$C_{S}$ $2\tau_{S}/\bigl{(}\rho U_{S}^{2}\bigr{)}$ Response
Friction coefficient associated with shear forcing
$C_{L}$ $2\tau_{L}/\bigl{(}\rho U_{L}^{2}\bigr{)}$ Response
Friction coefficient associated with LSC rolls for sheared RB
$C_{R}$ $2\tau_{R}/\bigl{(}\rho U_{R}^{2}\bigr{)}\equiv C_{L}(Re_{S}=0)$ Response
Friction coefficient associated with LSC rolls for pure RB
$C_{T}$ $2\tau_{T}/\bigl{(}\rho U_{T}^{2}\bigr{)}$ Response
Friction coefficient computed using the total velocity
$Nu$ $u_{z}\theta-\kappa\partial_{z}\theta\right>_{A,t}/\bigl{(}\kappa\Delta H^{-1}\bigr{)}$ Response
Nusselt number
@@
Using the exact relations for the global kinetic and thermal dissipation rates, Grossmann & Lohse (2000, 2001, 2002, 2004) offered a unifying theory for RB convection (hereafter referred to as the GL-theory) which, for any given $Ra$ and $Pr$, reliably predicts the value of $Nu$ and $Re_{L}$ (see also Stevens et al. (2013); Lohse & Shishkina (2023)). Presently, the theory has been applied to RB convection without imposed shear. The objective of this work is to extend the theoretical approach to sheared thermal convection. Scagliarini et al. (2014, 2015); Pirozzoli et al. (2017); Blass et al. (2020, 2021) have made progress in understanding the variation of heat transfer in RB convection with imposed shear. Scagliarini et al. (2014, 2015) proposed a model based on the concept of eddy viscosity and eddy diffusivity to explain the counter-intuitive initial decrease and subsequent increase in $Nu$ with increasing $Re_{b}$ for the PRB system. Blass et al. (2020) observed a similar effect in the CRB system and attributed it to the initial destruction of the large-scale flow organisation and the subsequent formation of large meandering flow structures (Hutchins & Marusic, 2007). They divided the flow into a buoyancy-dominated, a transitional, and a shear-dominated regime, based on the Monin–Obukhov length scale. Blass et al. (2021) further investigated the effect of $Pr$ on the variation of $Nu$ with $Re_{w}$ and concluded that the non-monotonic behaviour of $Nu(Re_{w})$ is a consequence of flow layering, plume sweeping, and bulk heat entrapment. Building on these findings, in this paper we propose a more general formulation applicable to sheared thermal convection, in the spirit of the GL-theory.
@@ 2.2 Kinetic energy, large scale circulation, and boundary layer thickness
An interpretation of the above equation (10) is that the velocity associated with the LSC is preferentially oriented along the direction orthogonal to the shear forcing, consistent with the flow structures observed by Pirozzoli et al. (2017); Blass et al. (2020).
This can be better understood by considering many LSC rolls each orientated at an angle $\alpha$ with the streamwise direction and studying the probability distribution of $\alpha$ given by $\phi(\alpha)$. Since the total velocity $U_{T}$ arises from a vector addition of the shear velocity and the LSC velocities, we can express it in terms of $\phi(\alpha)$ as
@@ 2.3 Estimating bulk and boundary layer contributions
which is, in turn, a sum of contributions from the LSC and the applied shear forcing. Note that the shear contribution here exactly matches that in the global relation (7). Next, we estimate $\epsilon_{u,bulk}$ by assuming that the velocity scale $U_{L}$ associated with the LSC is responsible for stirring the bulk with a kinetic energy that scales with $U_{L}^{2}$. However, the timescale of the stirring process is governed not by the velocity $U_{L}$ associated with the LSC but by $U_{T}$ which is the total velocity scale. As we shall find later in §3.2, this is a key assumption that explains the trend of $Nu$ with increasing $Re_{S}$ in the buoyancy-dominated regime. Following these assumptions, we write
@@
$\epsilon_{\theta,bulk}\sim\Delta^{2}\frac{U_{L}}{H}\frac{\lambda_{\theta}}{\lambda_{u}}\sim\kappa\frac{\Delta^{2}}{H^{2}}PrC_{L}Re_{L}^{2}Nu^{-1}\equiv\kappa\frac{\Delta^{2}}{H^{2}}\frac{Pr}{C_{L}}\frac{\left(C_{T}Re_{T}\right)^{2}}{Nu}$ (22)
@@ 2.4 Limiting regimes
Table 2: Scaling relations for the Nusselt number $Nu$ and LSC Reynolds number $Re_{L}$ in the buoyancy-dominated regime of sheared Rayleigh–Bénard convection. The first column indicates the GL regime, the second column indicates the bulk or boundary layer (BL) dominance of the kinetic dissipation rate with the applicable scaling estimate in the parenthesis, and the third column indicates the bulk or boundary layer (BL) dominance of the thermal dissipation rate with the applicable scaling estimate in the parenthesis. The fourth column indicates whether the kinetic boundary layer is thicker than the thermal boundary layer or vice-versa. The fourth and fifth columns indicate the scaling relations for $Nu/Nu_{R}$ and $Re_{L}/Re_{R}$ using the values of $Nu$ and $Re_{R}$ estimated for the pure Rayleigh–Bénard system from GL-theory.
@@
Using these approximations, we can provide scaling relations between $Nu/Nu_{R}$ and $Re/Re_{R}$ with $Nu_{R}$ and $Re_{R}$ being the Nusselt number and Reynolds number associated with the LSC for the pure RB system. Following the GL-theory, there are various regimes that can be relevant depending on whether the dissipation rates are dominated by boundary layer or bulk contributions, and whether the thermal boundary layer is thicker than the kinetic boundary layer. For each of these regimes, we combine (12) and (28) with the relevant estimates for the dominant dissipation rate contribution to give expressions for $Nu(Ra,Pr,C_{L},Re_{L},Re_{T})$ and $Re_{L}(Ra,Pr,C_{L},Re_{L},Re_{T})$. Since our approach is consistent with the GL-theory, the $Ra$ and $Pr$ dependence simply recovers the scaling relations $Nu_{R}(Ra,Pr)$ and $Re_{R}(Ra,Pr)$ found for the various regimes of pure RB, providing us with scaling relations for $Nu/Nu_{R}$ and $Re_{L}/Re_{R}$ that only depend on $C_{L}$, $Re_{L}$, and $Re_{T}$. In table 2, we outline the relevant estimates for the dissipation rates and present the resulting scaling relations.
In the buoyancy-dominated classical GL regimes, we can consider the Prandtl–Blasius scaling (23) to hold for small shear forcing (i.e., $Re_{S}\lesssim Re_{L}$). With this assumption, $C_{L}^{2}Re_{L}=1$, so the values of $Nu$ and $Re_{L}$ remain unchanged for buoyancy-dominated regimes $I$ and $III$, whereas for buoyancy-dominated regimes $II$ and $IV$, the non-monotonic behaviour of $Nu$ with increasing $Re_{S}$ becomes apparent. Although $Nu$ seems to decrease with increasing $Re_{S}$ in the buoyancy-dominant $II$ and $IV$ regimes, it is important to note that this behaviour is subject to the condition that the boundary layer is a laminar one of the Prandtl–Blasius type. If the boundary layer becomes turbulent, the expected decrease in $Nu$ in the buoyancy-dominated regime might disappear. In this study, we will explore the $Nu$ response in the buoyancy-dominated $II_{u}$ regime where we can still observe the decrease in $Nu$ with increasing $Re_{S}$ within reasonable computational cost.
@@ 3.1 Scheme and procedure
Figure 2: (a) The RMS horizontal velocity profile $u_{R}(z)$ normalized with the free fall velocity for RB system with $Ra=10^{7}$, $Pr=1.0$ is shown as a function of wall-normal height $z$, indicating the maximum value $U_{\mathrm{max}}$ occurring at a wall-normal height $\delta$. (b) $Re_{\mathrm{max}}$ obtained from $U_{\mathrm{max}}$ is plotted against the globally averaged RMS velocity $Re_{R}$ of the RB system. The dashed grey line indicates $Re_{\mathrm{max}}=Re_{RMS}$, showing that these two estimates are virtually identical. The values of $Ra$ are shown on the top for reference.
@@ 3.2 Global response parameters
Figure 3: (a) $Nu/Pr^{1/2}$ plotted against $Re_{S}$. The black solid line indicates (34), the dashed lines indicate (25), and the dotted lines indicate (32). (b) $Nu/Nu_{R}$ plotted against $Re_{S}/Re_{R}$. Dotted lines indicate (32). (c) $C_{S}$ plotted against $Re_{S}$. The black solid line indicates (33) and coloured dashed lines indicate (24). (d) $C_{S}$ normalised with $C_{R}$ plotted against $Re_{S}/Re_{R}$. Coloured dashed lines indicate (33) and black dashed line indicates (24). The data for PRB are indicated with plus markers and the data for CRB are indicated with dot markers.
@@
These equations show good agreement with the numerical data plotted in figures 3a-3d for the buoyancy-dominated regime.
Figure 4: (a) $Re_{T}/Re_{R}$ plotted against $Re_{S}/Re_{R}$. The black solid line indicates the relation (10). (b) $\left(\sqrt{Re_{T}^{2}-Re_{S}^{2}}\right)/Re_{R}$ plotted against $\sqrt{C_{S}}Re_{S}/Re_{R}$. The black dashed line indicates the relation (9) with $\gamma\approx 10.24$. The data for PRB are indicated with plus markers and the data for CRB are indicated with dot markers.

### v2

@@ 1 Introduction
The interplay between buoyancy and shear in mixed thermal convection can be studied by either adding Couette–type forcing to the Rayleigh–Bénard (RB) system (Ahlers et al., 2009; Lohse & Xia, 2010; Chilla & Schumacher, 2012; Xia, 2013; Shishkina, 2021; Ahlers et al., 2022; Lohse & Shishkina, 2023) to obtain the Couette–RB (CRB) system (Deardorff, 1965; Ingersoll, 1966; Hathaway & Somerville, 1986; Domaradzki & Metcalfe, 1988; Solomon & Gollub, 1990; Shevkar et al., 2019; Blass et al., 2020, 2021), or by applying a Poiseuille–type forcing to obtain the Poiseuille–RB (PRB) system (Zonta & Soldati, 2014; Scagliarini et al., 2014, 2015; Pirozzoli et al., 2017). A schematic of the two systems is shown in figure 1.
@@
Parameter Definition Type Short description
$Ra$ $\bigl{(}\beta gH^{3}\Delta\bigr{)}/(\nu\kappa)$ Control Rayleigh number
$Pr$ $\nu/\kappa$ Control Prandtl number
$Re_{F}$ $U_{F}H/\nu\equiv\sqrt{Ra/Pr}$ Control Reynolds number associated with free-fall velocity
$Re_{w}$ $U_{w}H/\nu$ Control Wall Reynolds number for CRB system
$Re_{b}$ $U_{b}H/\nu$ Control Bulk Reynolds number for PRB system
$Re_{S}$ $U_{S}H/\nu$ Control Reynolds number associated with shear forcing
$Re_{L}$ $U_{L}H/\nu$ Response Reynolds number associated with LSC rolls for sheared RB
$Re_{R}$ $U_{R}H/\nu\equiv Re_{L}(Re_{S}=0)$ Response Reynolds number associated with LSC rolls for pure RB (“wind Reynolds number”)
$Re_{T}$ $U_{T}H/\nu$ Response Reynolds number computed using the total velocity
$C_{S}$ $2\tau_{S}/\bigl{(}\rho U_{S}^{2}\bigr{)}$ Response Friction coefficient associated with shear forcing
$C_{L}$ $2\tau_{L}/\bigl{(}\rho U_{L}^{2}\bigr{)}$ Response Friction coefficient associated with LSC rolls for sheared RB
$C_{R}$ $2\tau_{R}/\bigl{(}\rho U_{R}^{2}\bigr{)}\equiv C_{L}(Re_{S}=0)$ Response Friction coefficient associated with LSC rolls for pure RB
$C_{T}$ $2\tau_{T}/\bigl{(}\rho U_{T}^{2}\bigr{)}$ Response Friction coefficient computed using the total velocity
$Nu$ $u_{z}\theta-\kappa\partial_{z}\theta\right>_{A,t}/\bigl{(}\kappa\Delta H^{-1}\bigr{)}$ Response Nusselt number
@@
Using the exact relations for the global kinetic and thermal dissipation rates, Grossmann & Lohse (2000, 2001, 2002, 2004) offered a unifying theory for RB convection (hereafter referred to as the GL-theory). For cylinders of unit aspect ratio, Stevens et al. (2013) have demonstrated that fitting the GL-theory at four data points from Funfschilling et al. (2005) at $Ra=2.96\times 10^{7}$ and $Ra=1.92\times 10^{10}$ with $Pr=4.38$, from Xia et al. (2002) at $Ra=2.24\times 10^{8}$ with $Pr=554$ and from Kerr & Herring (2000) at $Ra=10^{7}$ with $Pr=0.07$ using four free parameters can predict $Nu$ within $4\%$ of experimental and numerical results in most of the parameter space given by $10^{4}\leq Ra\leq 10^{14}$ and $10^{-3}\leq Pr\leq 10^{2}$ with only two small ranges that exhibit a greater $10\%$ disagreement. In the same paper Stevens et al. (2013), a fit of similar quality is achieved also for an aspect ratio Gamma = 1/2, with slightly different prefactors. This work has further been extended by Ahlers et al. (2022) to include the effects of aspect ratio between $1/32$ and $32$. The GL-theory has also been extended to the ultimate regime (Grossmann & Lohse, 2011), where the heat transport is considerably enhanced, as the laminar-type boundary layers become turbulent due to a non-normal–nonlinear instability; for a detailed discussion see Roche (2020) and Lohse & Shishkina (2023).
Presently, the GL-theory has been applied to RB convection without imposed shear. The objective of this work is to extend the theoretical approach to sheared thermal convection. Scagliarini et al. (2014, 2015); Pirozzoli et al. (2017); Blass et al. (2020, 2021) have made progress in understanding the variation of heat transfer in RB convection with imposed shear. Scagliarini et al. (2014, 2015) proposed a model based on the concept of eddy viscosity and eddy diffusivity to explain the counter-intuitive initial decrease and subsequent increase in $Nu$ with increasing $Re_{b}$ for the PRB system. Blass et al. (2020) observed a similar effect in the CRB system and attributed it to the initial destruction of the large-scale flow organisation and the subsequent formation of large meandering flow structures (Hutchins & Marusic, 2007). They divided the flow into a buoyancy-dominated, a transitional, and a shear-dominated regime, based on the Monin–Obukhov length scale. Blass et al. (2021) further investigated the effect of $Pr$ on the variation of $Nu$ with $Re_{w}$ and concluded that the non-monotonic behaviour of $Nu(Re_{w})$ is a consequence of flow layering, plume sweeping, and bulk heat entrapment. Building on these findings, in this paper we propose a more general formulation applicable to sheared thermal convection, in the spirit of the GL-theory.
@@ 2.2 Kinetic energy, large scale circulation, and boundary layer thickness
An interpretation of the above equation (10) is that the velocity associated with the LSC is preferentially oriented along the direction orthogonal to the shear forcing, consistent with the flow structures observed by Pirozzoli et al. (2017); Blass et al. (2020). Note that the validity of equation (10) is limited to sufficiently low values of $Re_{S}$ wherein the contribution to spanwise shear stresses from shear forcing is negligible in comparison to the contribution from the LSC rolls. At high shear forcing, the spanwise shear stresses generated by velocity fluctuations arising purely from shear forcing may no longer be negligible, in which case equation (10) no longer holds.
In the buoyancy-dominated regime, relation (10) can be better understood by considering many LSC rolls each orientated at an angle $\alpha$ with the streamwise direction and studying the probability distribution of $\alpha$ given by $\phi(\alpha)$. Here, we mean LSC rolls in a broad sense, without addressing the exact details of the flow organisation at this stage. Empirical observations regarding flow organisation are reported in section §3.3. Since the total velocity $U_{T}$ arises from a vector addition of the shear velocity and the LSC velocities, we can express it in terms of $\phi(\alpha)$ as
@@ 2.3 Estimating bulk and boundary layer contributions
which is, in turn, a sum of contributions from the LSC and the applied shear forcing. Note that the shear contribution here exactly matches that in the global relation (7). Next, we estimate $\epsilon_{u,bulk}$. Here, it is important to note that the bulk dissipation rate is dominated by the contribution from LSC rolls, while the contribution from applied shear forcing is much smaller. Therefore, we only focus on estimating the contribution to $\epsilon_{u,bulk}$ from the LSC rolls by assume that the velocity scale $U_{L}$ associated with the LSC rolls is responsible for stirring the bulk with a kinetic energy that scales with $U_{L}^{2}$. However, LSC rolls are swept at the boundary layer height due to the applied shear forcing and the timescale of the stirring process is governed not by the velocity $U_{L}$ associated with the LSC but by $U_{T}$ which is the total velocity scale. As we shall find later in §3.2, this is a key assumption that explains the trend of $Nu$ with increasing $Re_{S}$ in the buoyancy-dominated regime. Following these assumptions, we write
@@
At high-shear forcing, the contribution of the boundary layer to the dissipation rate arising from shear forcing dominates the bulk contribution from LSC rolls. In the limiting case of passive transport in Couette/Poiseuille flow, one can rigorously derive that the total kinetic dissipation rate $\epsilon_{u}=(\nu^{3}/H^{4})C_{S}Re_{S}^{3}$. In this sense, the kinetic dissipation rate of sheared RB at high shear forcing will always be dominated by the boundary layer contribution and estimating $\epsilon_{u,bulk}$ in the shear-dominated regimes is redundant.
@@
$\epsilon_{\theta,bulk}\sim\Delta^{2}\frac{U_{L}}{H}\frac{\lambda_{\theta}}{\lambda_{u}}\sim\kappa\frac{\Delta^{2}}{H^{2}}PrC_{L}Re_{L}^{2}Nu^{-1}\equiv\kappa\frac{\Delta^{2}}{H^{2}}\frac{Pr}{C_{L}}\frac{\left(C_{T}Re_{T}\right)^{2}}{Nu}.$ (22)
We expect equations (19) - (22) to be valid in both buoyancy-dominated and shear-dominated regimes. This is accommodated by the fact that the dependencies $C_{T}(Re_{T})$ and $C_{L}(Re_{L})$ behave differently in the buoyancy-dominated and shear-dominated regimes. Compatibility of these relations in the limiting regimes is explained in the subsequent subsection §2.4.
@@ 2.4 Limiting regimes
Table 2: Scaling relations for the Nusselt number $Nu$ and LSC Reynolds number $Re_{L}$ in the buoyancy-dominated regime of sheared Rayleigh–Bénard convection with a laminar Prandtl–Blasius type boundary layer. The first column indicates the GL regime, the second column indicates the bulk or boundary layer (BL) dominance of the kinetic dissipation rate with the applicable scaling estimate in the parenthesis, and the third column indicates the bulk or boundary layer (BL) dominance of the thermal dissipation rate with the applicable scaling estimate in the parenthesis. The fourth column indicates whether the kinetic boundary layer is thicker than the thermal boundary layer or vice-versa. The fourth and fifth columns indicate the scaling relations for $Nu/Nu_{R}$ and $Re_{L}/Re_{R}$ using the values of $Nu$ and $Re_{R}$ estimated for the pure Rayleigh–Bénard system from GL-theory.
@@
Using these approximations, we can provide scaling relations between $Nu/Nu_{R}$ and $Re/Re_{R}$ for buoyancy-dominated sheared RB system with $Nu_{R}$ and $Re_{R}$ being the Nusselt number and Reynolds number associated with the LSC for the pure RB system. Following the GL-theory, there are various regimes that can be relevant depending on whether the dissipation rates are dominated by boundary layer or bulk contributions, and whether the thermal boundary layer is thicker than the kinetic boundary layer. For each of these regimes, we combine (12) and (28) with the relevant estimates for the dominant dissipation rate contribution to give expressions for $Nu(Ra,Pr,C_{L},Re_{L},Re_{T})$ and $Re_{L}(Ra,Pr,C_{L},Re_{L},Re_{T})$. Since our approach is consistent with the GL-theory, the $Ra$ and $Pr$ dependence simply recovers the scaling relations $Nu_{R}(Ra,Pr)$ and $Re_{R}(Ra,Pr)$ found for the various regimes of pure RB, providing us with scaling relations for $Nu/Nu_{R}$ and $Re_{L}/Re_{R}$ that only depend on $C_{L}$, $Re_{L}$, and $Re_{T}$. In table 2, we outline the relevant estimates for the dissipation rates and present the resulting scaling relations. Is it important to note that these scaling relations are only applicable to buoyancy-dominated sheared RB flows with scaling-wise laminar Prandtl–Blasius type boundary layers.
In the buoyancy-dominated classical GL regimes, we can consider the Prandtl–Blasius scaling (23) to hold for small shear forcing (i.e., $Re_{S}\lesssim Re_{L}$). With this assumption, $C_{L}^{2}Re_{L}\approx 1$, so the values of $Nu$ and $Re_{L}$ remain unchanged for buoyancy-dominated regimes $I$ and $III$, whereas for buoyancy-dominated regimes $II$ and $IV$, the non-monotonic behaviour of $Nu$ with increasing $Re_{S}$ becomes apparent. Although $Nu$ seems to decrease with increasing $Re_{S}$ in the buoyancy-dominant $II$ and $IV$ regimes, it is important to note that this behaviour is subject to the condition that the boundary layer is a laminar one of the Prandtl–Blasius type. If the boundary layer becomes turbulent, the expected decrease in $Nu$ in the buoyancy-dominated regime might disappear. In this study, we will explore the $Nu$ response in the buoyancy-dominated $II_{u}$ regime where we can still observe the decrease in $Nu$ with increasing $Re_{S}$ within reasonable computational cost.
@@ 3.1 Scheme and procedure
Figure 2: (a) The RMS horizontal velocity profile $u_{R}(z)$ normalized with the free fall velocity for RB system with $Ra=10^{7}$, $Pr=1.0$ is shown as a function of wall-normal height $z$, indicating the maximum value $U_{\mathrm{max}}$ occurring at a wall-normal height $\delta$. (b) $Re_{\mathrm{max}}$ obtained from $U_{\mathrm{max}}$ is plotted against the globally averaged RMS velocity $Re_{R}$ of the RB system. The dashed grey line indicates $Re_{\mathrm{max}}=Re_{RMS}$, showing that these two estimates are virtually identical. The values of $Ra$ are shown on the top for reference. This figure is also available as an interactive JFM Notebook.
@@ 3.2 Global response parameters
Figure 3: (a) $Nu/Pr^{1/2}$ plotted against $Re_{S}$. The black solid line indicates (34), the dashed lines indicate (25), and the dotted lines indicate (32). (b) $Nu/Nu_{R}$ plotted against $Re_{S}/Re_{R}$. Dotted lines indicate (32). (c) $C_{S}$ plotted against $Re_{S}$. The black solid line indicates (33) and coloured dashed lines indicate (24). (d) $C_{S}$ normalised with $C_{R}$ plotted against $Re_{S}/Re_{R}$. Coloured dashed lines indicate (33) and black dashed line indicates (24). The data for PRB are indicated with plus markers and the data for CRB are indicated with dot markers. This figure is also available as an interactive JFM Notebook.
@@
These equations show good agreement with the numerical data plotted in figures 3a-3d for the buoyancy-dominated regime. Note that equations 3.1 and 3.3 do not explicitly state the dependence of $Nu_{R}(Ra,Pr)$ or $Re_{R}(Ra,Pr)$. There is no pure scaling exponent for $Nu_{R}(Ra)$ in pure RB for regime $II_{u}$. The values of $Nu_{R}$ and $Re_{R}$ used for figures 3b and 3d are obtained from numerical simulations of pure RB. The present entension to the GL-theory assumes that the values of $Nu_{R}$ and $Re_{R}$ are known a-priori, and only attemts to provide scaling relations for the normalised quantities $Nu/Nu_{R}(Re_{S}/Re_{R})$ and $C_{S}/C_{R}(Re_{S}/Re_{R})$ in the buoyancy-dominated regime.
Figure 4: (a) $Re_{T}/Re_{R}$ plotted against $Re_{S}/Re_{R}$. The black solid line indicates the relation (10). (b) $\left(\sqrt{Re_{T}^{2}-Re_{S}^{2}}\right)/Re_{R}$ plotted against $\sqrt{C_{S}}Re_{S}/Re_{R}$. The black dashed line indicates the relation (9) with $\gamma\approx 10.24$. The data for PRB are indicated with plus markers and the data for CRB are indicated with dot markers. This figure is also available as an interactive JFM Notebook.